### v1

@@ Abstract
Unmanned Aerial Vehicles (UAVs) have become essential tools in inspection and emergency response operations due to their high maneuverability and ability to access hard-to-reach areas. However, their limited battery life significantly restricts their use in long-duration missions. This paper presents a novel tethered marsupial robotic system composed of a UAV and an Unmanned Ground Vehicle (UGV), specifically designed for autonomous, long-duration inspection tasks in Global Navigation Satellite System (GNSS)-denied environments. The system extends the UAV’s operational time by supplying power through a tether connected to high-capacity battery packs carried by the UGV. We detail the hardware architecture based on off-the-shelf components to ensure replicability and describe our full-stack software framework, which is composed of open-source components and built upon the Robot Operating System (ROS). The proposed software architecture enables precise localization using a Direct LiDAR Localization (DLL) method and ensures safe path planning and coordinated trajectory tracking for the integrated UGV–tether–UAV system. We validate the system through three field experiments: (1) a manual flight endurance test to estimate the operational duration, (2) an autonomous navigation test, and (3) an inspection mission to demonstrate autonomous inspection capabilities. Experimental results confirm the robustness and autonomy of the system, its capacity to operate in GNSS-denied environments, and its potential for long-endurance, autonomous inspection and monitoring tasks.
keywords:
GNSS-denied areas, Tethered robots, Long-duration inspection, Marsupial system
@@ 1 Introduction
The research and deployment of Unmanned Aerial Vehicles (UAVs) have experienced exponential growth during the twenty-first century in a broad variety of civil applications [1]. Some applications which include UAVs as the main agents are: road traffic monitoring, remote sensing, search and rescue operations in emergency situations, security and surveillance over an area, assessment and spraying in agriculture [2], forest firefighting [3, 4]. In them, UAVs are widely used due to their high maneuverability, mobility, and the decreasing costs of both purchase and maintenance. Other applications that have benefited from the use of UAVs are related to inspection tasks to analyze and monitor civil infrastructure [5] [6]. These tasks use the capabilities of the UAV to move within closed, dangerous, or risky environments, or to reach objectives that are at a great distance. For example, UAVs have been used in structure assessment on buildings and bridges [7], power line inspection, facade and indoor fire monitoring and extinguishing [4, 8] and even sewer inspection [9].
@@
However, a significant limitation of small and medium-sized UAVs is their flight duration, which hinders their ability to conduct comprehensive inspection tasks. In fact, the majority of commercial platforms exhibit flight times of less than thirty minutes, typically measured under laboratory conditions and without any additional payload, such as computing devices or sensors.
@@
There are some alternatives developed for extending the flight duration of UAVs. The most common approach is to perform battery swapping either manually, thus requiring human intervention, or automatically with battery swapping stations [10], requiring the installation of additional infrastructure. Another alternative is to use a swarm of drones [11], which greatly increases the complexity of the system and its maintenance costs. Finally, the use of a tethered UAV fixed to a base station that supplies its power [12] is another approach more related to the one proposed in this paper. In fact, tethered UAVs have gathered the attention of the community [13] with applications as diverse as collaborative load transportation [14], high bandwidth communications, [15, 16] and airborne wind energy [17]. Regrettably, the use of a fixed tether at a given location significantly restricts the operational range of the UAV, thereby preventing it from safely accessing certain areas of interest.
@@
In this paper, we present an autonomous tethered marsupial robotic configuration for long-duration inspection tasks. Our system utilizes an Unmanned Ground Vehicle (UGV) to transport battery packs that supply power through a cable to a UAV equipped with the sensors needed to carry out the inspection. This approach is based on a philosophy of robotic collaboration, offering remarkable advantages over traditional standalone systems by leveraging the unique strengths of each robotic agent [18]. A case of success of a marsupial robotic system can be found in the DARPA 2021 Subterranean Challenge [19], where the CSIRO Data61 Team reached second place [20]. The team used a marsupial system in which a UGV transports a UAV to a take-off point to carry out inspection tasks, which reduces the UAV energy consumption. Another successful case is the Mars 2020 rover and helicopter [21] which is being used for planetary exploration on Mars. In this system, the helicopter augments the capabilities of the rover, exploring large areas faster than the rover, providing reconnaissance on target locations and safe-to-traverse routes. In both applications, the concept of collaboration based on the robotic strengths is used; however, the robots are not tethered to provide energy, and the UAV operation time is still limited by the power it can carry.
@@
An important milestone in the context of marsupial tethered robots is introduced in [22], where the authors describe the development of a tethered UAV system operating alongside a UGV for inspecting stone-mine pillars and assessing structural integrity in underground mines. The UAV is powered via a tether from the UGV, allowing for extended operation without relying solely on its onboard battery. The UGV carries additional batteries, providing continuous power through the tether and effectively addressing the UAV’s flight-time limitations. Despite the robustness of the implemented system, the work focuses on autonomous navigation and inspection around stone pillars, rather than enabling navigation and inspection in general or more complex environments. Furthermore, the article does not provide an analysis of the UAV’s long-term inspection capabilities. A similar collaborative UAV-UGV system connected by a power tether is presented in [12]. This system leverages the strengths of both platforms to extend autonomous navigation in partially mapped environments: the UGV, equipped with a large power supply, powers the UAV, while the UAV contributes aerial surveillance, enhanced situational awareness, and obstacle detection for the UGV. However, this work lacks a detailed description of the system architecture and does not analyze flight duration.
@@
Another approach based on a marsupial system is presented in [23], where the authors introduce an optimal path-planning strategy for a marsupial robotic system consisting of a UGV, a UAV, and a tether connecting both robots. The system follows a sequential navigation strategy through 3D environments, avoiding both ground and aerial obstacles to ensure a collision-free trajectory for the team. Nonetheless, this work does not present a complete analysis of the system’s suitability for autonomous inspection tasks or evaluate its flight endurance. In [24], the authors introduce a collaborative UGV-UAV system for reliable localization and autonomous operation in Global Navigation Satellite System (GNSS)-denied environments. The UGV, equipped with a 3D LiDAR, a fisheye camera, an Ultra-WideBand (UWB) radio, and an Inertal Measurement Unit (IMU), handles mapping and localization, while the UAV uses shared IMU and laser altimeter data from the UGV. An Error-State Extended Kalman Filter fuses these sensor data to estimate the UAV’s relative position, enabling real-time tracking and 3D mapping. However, the system does not consider the tether in its planning process, nor does it address UAV flight-time limitations. In [25], a simulator based on the Robotic Operating System (ROS) version 2 and the Gazebo simulator is presented for marsupial systems where a UAV and a UGV are connected by an adjustable-length hanging tether. This configuration enables realistic modeling of the interactions between the robots and the winch that regulates the tether. Its primary contribution is to offer a tool for validating and testing planning and control algorithms, thereby reducing the costs and risks of physical experiments. However, the work is limited to serving as a tool for testing navigation modules, and it does not facilitate field experiments to quantify the viability of this type of system, the UAV’s flight duration, or even the battery charge status.
@@
This paper extends our previous work presented in [26], in which we presented a short description of the proposed system for long-duration missions. As the extension, this paper gives a detailed description of the main onboard software modules used for localization and navigation purposes. In addition, we present a larger set of validation experiments, including quantitative localization and navigation results. Thus, the main contributions of this paper are as follows:
@@
    We present our marsupial robotic system’s hardware and software design (see Fig. 1), which considers a UGV tied to a UAV, based on off-the-shelf components in contrast to existing approaches in the literature. In our system, we use a tether that connects the UAV to the UGV. This tether provides the UAV with power, enabling long-term inspection missions.
@@
    We present the architecture of our software solution. The different software modules, developed in-house, are released as open-source. All the modules have been incrementally validated to reach the final goal of performing autonomous inspection missions with a tethered UAV-UGV marsupial system. This is a significant contribution w.r.t. [26, 27, 28].
@@
    We pay special emphasis on describing our internally developed localization and navigation systems. In particular, the localization system is based on the one presented in our previous work [27], but is adapted for the current case and for providing multi-robot localization in Section 3. Regarding navigation, we present a reformulation of our trajectory planner from [28], adapted to account for the constraints of the proposed marsupial system presented in this paper.
The contributions of the paper are validated in several field experiments, including an inspection experiment of a civil infrastructure.
The paper is structured as follows: Section 2 provides a basic description of the hardware of the UGV, the UAV and the tether systems, including details on the power system for the UAV. It also presents the basic software architecture of the system. Then, we describe in detail the localization and navigation systems used in autonomous inspection missions in Sections 3 and 4, respectively. Our marsupial system has been extensively validated in experimental tests on three different scenarios. These results are presented in Section 5. Later, we present an extended discussion of our results in Section 6. Finally, the conclusions and future research directions are detailed in Section 7.
Figure 1: Our marsupial system performing autonomous inspection tasks in two buildings with structural damage in the University Pablo de Olavide, Seville (Spain). (a) Detail of the UAV at the abandoned thermal station (Scenario 2). (b) Tethered marsupial system at the old theater building (Scenario 3).
2 System Overview
In this section, we describe the main parts of our marsupial system. The system consists of a UGV, a UAV, and a tether system (see Fig. 2), which provides the UAV with power from additional battery packs onboard the UGV. In this way, we can greatly extend the flight duration of the UAV. The UAV, the UGV and the tether system are described in Sections 2.1, 2.2, and 2.3, respectively. Section 2.4 analyzes the power requirements to calculate the capacity of the batteries on the UGV. Finally, Section 2.5 gives an overview of our software solution for autonomous operation.
@@ 2 System Overview
Figure 2: Marsupial System, UAV tied to UGV. (a) Detail of the UGV equipped with the auxiliary structure that stores the systems of the marsupial configuration. (b) Complete system in an initial position before the UAV takeoff.
2.1 Unmanned Ground Vehicle - UGV
We use an ARCO ground platform from IDMind [29], which is a 4-wheeled holonomic robot with an independent traction system designed for the delivery of heavy loads in factories (see Fig. 3). The platform carries its own batteries, which are two LiFePO4 24V, for motors and electronics. ARCO has an autonomy of 3 hours of operation with a maximum payload of over $100\ kg$ and a maximum speed of 0.8 m/s.
@@ 2.1 Unmanned Ground Vehicle - UGV
To carry out the task, we have properly equipped ARCO with the necessary components to achieve the marsupial configuration (see Fig. 2a), which are detailed below:
  • Onboard Computer, model MSI CubiN 8GL-001BEU with an Intel i5 processor.
  • A Sparkfun Razor 9 DoF IMU placed inside of the platform. Used for mapping and localization.
  • LiDAR: A 3D OS1-16 Ouster LiDAR, located on top of the robot. It generates 320k points per second, ranging from 0.2 to 120 meters. This information is used for localization, mapping and obstacle avoidance.
  • Auxiliary structure. Designed primarily to position the tether system at the top of the UGV. It is built from Boch RexRoth aluminum profiles.
  • Battery pack. This is part of the tether power system. Two LiFePO4 12V batteries store the energy that is provided to the UAV to increase its flight time.
  • Inverter. This is part of the tethered power system. It converts the DC energy from the batteries into AC energy for the tether system.
  • Reel System. It is placed on top of the auxiliary structure. It connects the Power System with the UAV.
  • Router. The UGV is equipped with a NIGHTHAWK NetGear AX5400 Router. In our setup, UGV has a wired connection via Ethernet, while the Base Station and the UAV connect via 5G WiFi.
2.2 Unmanned Aerial Vehicle - UAV
Regarding the aerial platform, the model used for the experiments is the DJI Matrice 210 V2 (M210) (see Fig. 4). We selected this platform because of the robustness of the hardware and software, the support that the provider offers, and its ROS integration by means of a Software Development Kit (SDK).
The drone’s structure was adapted in order to carry an i7 NUC computer and a LiDAR, as shown in Fig. 4. The main sensors and systems on the UAV are:
  • M210 onboard sensors. Our UAV includes a single GNSS receiver, an IMU and a barometric altimeter.
  • 3D LiDAR sensor. An OS1-16 Ouster LiDAR is used for environment mapping, localization, and odometry.
  • USB WiFi access point. A D-Link AX1800 Wi-Fi USB Adapter is used to increase the robustness, range and performance of the WiFi connection at the UAV.
@@ 2.3 Tether System
We use the system LIGH-T V4 Tethered Station (see Fig. 5), provided by the company Elistair [30]. We selected this system due to the support that the company provides and its compatibility with the M210, among other DJI models. The cable of the system ends in an adapter, which is located in the drone battery compartment to provide energy to the drone. Note that a fully charged battery has to be equipped in the UAV in the remaining battery slot on the M210. This battery acts as a backup that is used in case of a disconnection between the UAV and the UGV.
The maximum length of the tether is around seventy meters. Besides, the system works with an AC input. Thus, an inverter is required to connect it to the battery pack. Finally, the reel is connected to a DC motor, which can perform a constant tension to the cable ranging from one to three Newtons.
Figure 5: (a) LIGH-T V4 Tethered Station, including switches, a reel for the cable, and the battery connector. (b) Detail of the battery connector.

### v2

@@ Abstract
Unmanned Aerial Vehicles (UAVs) have become essential tools in inspection and emergency response operations due to their high maneuverability and ability to access hard-to-reach areas. However, their limited battery life significantly restricts their use in long-duration missions. This paper presents a tethered marsupial robotic system composed of a UAV and an Unmanned Ground Vehicle (UGV), specifically designed for autonomous, long-duration inspection tasks in Global Navigation Satellite System (GNSS)-denied environments. The system extends the UAV’s operational time by supplying power through a tether connected to high-capacity battery packs carried by the UGV. Our work details the hardware architecture based on off-the-shelf components to ensure replicability and describes our full-stack software framework used by the system, which is composed of open-source components and built upon the Robot Operating System (ROS). The proposed software architecture enables precise localization using a Direct LiDAR Localization (DLL) method and ensures safe path planning and coordinated trajectory tracking for the integrated UGV–tether–UAV system. We validate the system through three sets of field experiments involving (i) three manual flight endurance tests to estimate the operational duration, (ii) three experiments for validating the localization and the trajectory tracking systems, and (iii) three executions of an inspection mission to demonstrate autonomous inspection capabilities. The results of the experiments confirm the robustness and autonomy of the system in GNSS-denied environments. Finally, all experimental data have been made publicly available to support reproducibility and to serve as a common open dataset for benchmarking.
@@ 1 Introduction
Research and deployment of Unmanned Aerial Vehicles (UAVs) have experienced exponential growth during the 21st century in a broad variety of civil applications [1]. Some applications that include UAVs as the main agents are road traffic monitoring, remote sensing, search and rescue operations in emergency situations, security and surveillance over an area, assessment and spraying in agriculture [2], forest firefighting [3]. In them, UAVs are widely used due to their high maneuverability, mobility, and the decreasing costs of both purchase and maintenance. Other applications that have benefited from the use of UAVs are related to inspection tasks to analyze and monitor civil infrastructure [4, 5]. These tasks use the capabilities of the UAV to move within closed, dangerous, or risky environments or to reach objectives that are at a great distance. For example, UAVs have been used in the evaluation of the structure of buildings and bridges [6], the inspection of power lines, the monitoring and extinguishing of fires in the facade and interior of buildings [3].
@@
However, a significant limitation of small and medium-sized UAVs is their flight duration, which hinders their ability to conduct comprehensive inspection tasks. In fact, the majority of commercial platforms exhibit flight times of less than thirty minutes, typically measured under laboratory conditions and without any additional payload, such as computing devices or sensors [7]. Fully equipped autonomous drones carrying the necessary sensor payload seldom achieve flight times of more than twenty minutes [8].
@@
There are some alternatives to develop long-duration inspection tasks. The most common approach is to perform battery swapping either manually, thus requiring human intervention, or automatically with battery swapping stations [9], requiring the installation of additional infrastructure. Another alternative is to use a team of robots flying either alternately or simultaneously, so that the inspection mission can be carried out without interruptions or in a shorter amount of time [10]. However, the use of a team of UAVs greatly increases the complexity of the system and its acquisition and maintenance costs. Finally, tethered UAVs connected to a base station that supplies them with power have attracted the attention of the research community [11, 12] with applications as diverse as collaborative load transportation [13], high-bandwidth communications [14, 15] and airborne wind energy [16]. Regrettably, the use of a fixed tether at a specific location significantly limits the operational range of the UAV, thereby preventing safe access to certain areas of interest.
@@
In this paper, we present an autonomous tethered marsupial robotic configuration for long-duration inspection tasks. Our system utilizes an Unmanned Ground Vehicle (UGV) to transport battery packs that supply power through a cable to a UAV equipped with the sensors needed to carry out the inspection. This approach is based on a philosophy of robotic collaboration, offering remarkable advantages over traditional standalone systems by leveraging the unique strengths of each robotic agent [17]. A case of success of a marsupial robotic system can be found in the Defense Advanced Research Projects Agency (DARPA) 2021 Subterranean Challenge [18], where the CSIRO Data61 Team reached second place [19]. The team used a marsupial system in which a UGV transports a UAV to a take-off point to carry out inspection tasks, which reduces the UAV’s energy consumption. Another successful case is the Mars 2020 rover and helicopter [20] which are being used for planetary exploration on Mars. In this system, the helicopter augments the capabilities of the rover, exploring large areas faster than the rover and providing reconnaissance on target locations and safe-to-traverse routes. In both applications, the concept of collaboration based on the robotic strengths is used; however, the robots are not tethered to provide energy, so the UAV operation time is still limited by the power it can carry.
@@
Following the tethered marsupial approach, an important milestone is introduced in [21], where the authors describe the development of a tethered UAV system operating alongside a UGV for inspecting stone-mine pillars and assessing structural integrity in underground mines. The UAV is powered via a tether from the UGV, allowing for extended operation without relying solely on its onboard battery. The UGV carries additional batteries, providing continuous power through the tether and effectively addressing the UAV’s flight-time limitations. Despite the robustness of the implemented system, the work focuses on autonomous navigation and inspection around stone pillars, rather than enabling navigation and inspection in general or more complex environments. Furthermore, the article does not provide an analysis of the UAV’s long-term inspection capabilities. A similar collaborative UAV-UGV system connected by a power tether is presented in [22]. This system leverages the strengths of both platforms to extend autonomous navigation in partially mapped environments: the UGV, equipped with a large power supply, powers the UAV, while the UAV contributes aerial surveillance, enhanced situational awareness, and obstacle detection for the UGV. However, this work lacks a detailed description of the system architecture and does not analyze flight duration.
@@
Another approach based on a marsupial system is presented in [23], where the authors introduce an optimal path-planning strategy for a marsupial robotic system consisting of a UGV, a UAV, and a tether connecting both robots. The system follows a sequential navigation strategy through 3D environments, avoiding both ground and aerial obstacles to ensure a collision-free trajectory for the team. Nonetheless, this work does not present a complete analysis of the system’s suitability for autonomous inspection tasks or evaluate its flight endurance. In [24], the authors introduce a collaborative UGV-UAV system for reliable localization and autonomous operation in Global Navigation Satellite System (GNSS)-denied environments. The UGV, equipped with a 3D LiDAR, a fisheye camera, an Ultra-WideBand (UWB) radio, and an Inertial Measurement Unit (IMU), handles mapping and localization, while the UAV uses shared IMU and laser altimeter data from the UGV. An Error-State Extended Kalman Filter fuses this sensor data to estimate the UAV’s relative position, enabling real-time tracking and 3D mapping. However, the system does not consider the tether in its planning process, nor does it address UAV flight-time limitations. In [25], a simulator based on the Robotic Operating System (ROS) version 2 and the Gazebo simulator is presented for marsupial systems where a UAV and a UGV are connected by an adjustable-length hanging tether. This configuration enables realistic modeling of the interactions between the robots and the winch that regulates the length of the tether. Its primary contribution is to offer a tool for validating and testing planning and control algorithms, thereby reducing the costs and risks of physical experiments. However, the work is limited to serving as a tool for testing navigation modules, and it does not facilitate field experiments to quantify the viability of this type of system, the UAV’s flight duration, or even the battery charge status.
@@
This paper extends our previous work presented in [26], in which we presented a short description of the proposed system for long-duration missions. As an extension, this paper gives a detailed description of the main onboard software modules used for control, localization and navigation purposes. Moreover, we present here extended field reports that show that our proposed system can perform autonomous inspection tasks in cluttered environments, which are still a challenge in tethered UAV-UGV systems [12]. To this end, we present a set of validation experiments, including quantitative localization and navigation results. Thus, the main contributions of this paper are as follows:
@@
    We present the hardware and software design of the proposed marsupial robotic system as a whole (see Figure 1). It consists of a UGV linked to a UAV and is built using off-the-shelf components in contrast to existing approaches in the literature [21]. In our system, we use a cable that connects the UAV to the UGV. This tether supplies the UAV with power, enabling long-term inspection missions. To the best of our knowledge, this is the first system to demonstrate a continuous operation exceeding two hours.
@@
    We present the architecture of our software solution. The different software modules, developed in-house, are released as open-source. All the modules have been incrementally validated to reach the final goal of performing autonomous inspection missions with a tethered UAV-UGV marsupial system. This is a significant contribution with respect to [26, 27, 28].
@@
    We pay special emphasis on describing our developed localization and navigation systems. In particular, the localization system is based on the one presented in our previous work [27], but is adapted for the current case and for providing multi-robot localization in Section 3. Regarding navigation, we present a reformulation of our trajectory planner from [28], adapted to account for the constraints of the proposed marsupial system. In particular, in this paper we always consider a straight tether instead of the variable-length hanging tether approach presented in [28].
The paper is structured as follows: Section 2 provides a basic description of the hardware of the UGV, the UAV and the tether systems, including details on the power system for the UAV. It also presents the basic software architecture of the system. Then, we describe in detail the localization and navigation systems used in autonomous inspection missions in Sections 3 and 4, respectively. Our marsupial system has been extensively validated in experimental tests in three different scenarios. These results are presented in Section 5. Later, we present an extended discussion of our results in Section 6. Finally, the conclusions and future research directions are detailed in Section 7.
@@
Figure 1: Our marsupial system performing autonomous inspection tasks in two buildings with structural damage in the University Pablo de Olavide (UPO), Seville (Spain). (a) Detail of the UAV at the abandoned thermal station (Scenario 2). (b) Tethered marsupial system at the old theater building (Scenario 3).
2 Materials and Methods
In this section, we describe the main components of our marsupial system. The system consists of a UGV, a UAV, and a tether system (see Figure 2), which supplies the UAV with power from additional battery packs onboard the UGV. This configuration allows us to significantly extend the flight duration of the UAV. The UGV, the UAV and the tether system are described in Sections 2.1, 2.2 and 2.3, respectively. Section 2.4 analyzes the power requirements used to determine the capacity of the batteries on the UGV. Finally, Section 2.5 gives an overview of our software solution for autonomous operation.
Refer to caption Refer to caption
(a) (b)
Figure 2: Marsupial system, UAV tied to UGV. (a) Detail of the UGV equipped with the auxiliary structure that stores the systems of the marsupial configuration. (b) Marsupial system in an initial position before the experiment.
2.1 Unmanned Ground Vehicle (UGV)
We use an ARCO ground platform from IDMind, which is a four-wheeled holonomic robot with an independent traction system designed for the delivery of heavy loads in factories (see Figure 3). The base platform has two batteries for motors and electronics, respectively, that give ARCO an autonomy of three hours of operation with a maximum payload of over one hundred kilograms and a maximum speed of 0.8 m/s.
@@ 2.1 Unmanned Ground Vehicle (UGV)
Additionally, we have properly equipped ARCO with the necessary components to achieve the marsupial configuration (see Figure 2a). Most importantly, the ARCO platform includes an internal computer that runs the internal software responsible for processing data from its IMU and 3D LiDAR sensors to localize the platform in real time. It also provides the rest of the system with high-speed 5G Wi-Fi connectivity via its integrated router. Finally, the tether system and additional batteries for powering the UAV are installed on the platform. All components of the ARCO platform are listed in Table 1.
Table 1: Main components of the ARCO UGV platform.
Component Model Quantity
Onboard Computer MSI Cubin 8GL-001BEU (i5-11400) 1
IMU Sparkfun Razor 9 DoF 1
LiDAR Ouster OS1-16 1
UGV Internal Battery LiFePO 75 Ah@24 V 2
UAV Power Battery Joiry LiFePO 150 Ah@12 V 2
Inverter Green Cell 2000 W/4000 W 1
Tether System Elistair LIGH-T V4 Tethered Station (LTS4) 1
WiFi Router NIGHTHAWK NetGear AX5400 1
2.2 Unmanned Aerial Vehicle (UAV)
Regarding the aerial platform, the model used for the experiments is the DJI Matrice 210 V2 (M210) (see Figure 4). We selected this platform because of the robustness of the hardware and software, the support that the provider offers, and its ROS integration by means of a Software Development Kit (SDK).
The main additions are its onboard computer and a LiDAR sensor, as shown in Figure 4. In addition, we have installed an external access point to increase the throughput and robustness of the WiFi connectivity (see Table 2).
@@ 2.2 Unmanned Aerial Vehicle (UAV)
Table 2: Main components added to the UAV platform
Component Model
Onboard computer NUC8i7BEH 32GB RAM
LiDAR Ouster OS1-16
Internal battery DJI TB55 (backup)
WiFi Access Point D-Link AX1800 Wi-Fi USB Adapter
@@ 2.3 Tether System
We use a LIGH-T V4 Tethered Station (LTS4), provided by the company Elistair [29] (see Figure 5) to power the UAV, due to its compatibility with the M210, among other DJI models. The cable of the system ends in an adapter, which is located in one of the two drone battery slots to provide energy to the drone. The maximum length of the cable is around seventy meters. Note that a fully charged battery has to be equipped in the UAV in the remaining battery slot on the M210. This battery acts as a backup that is used in case of a disconnection between the UAV and the LTS4. The cable is connected to a DC motor, which can perform a constant tension on the cable ranging from one to three Newtons.
Figure 5: (a) LIGH-T V4 Tethered Station (LTS4), including switches, a reel for the cable, and the battery connector. (b) Detail of the battery connector.